no.of faradays required to reduce 2.24l of 02 to H20 at STP if the current efficiency is 75%

Grade:12

1 Answers

AMIT AHIRE
5 years ago

From Faraday's electro-chemical laws, we know that one faraday of charge is required to liberate one equivalent of a substance.

As one mole of oxygen (or any gas, for that matter) occupies a volume of 22400cc at STP, the number of moles of oxygen gas are: 5600/22400=¼=0.25 moles of oxygen.

The gram equivalents of oxygen are obtained as moles×n factor=0.25×2=0.5 equivalents of oxygen gas.

Now from the first law, we can say that 0.5 faraday charge liberates 5600cc oxygen gas.

The answer is 0.5 Faraday charge.